# PAPI-Reg: Patch-to-Pixel Solution for Efficient Cross-Modal Registration between LiDAR Point Cloud and Camera Image

### v1

@@ 3.5 Loss Function
During training, we do not use top-k in the matching network, instead, we use the ground truth patch correspondences for subsequent pixel-level matching directly. For each patch correspondence, the network calculates $M$ assignment matrices $\mathbf{p}$. By dividing the actual correspondences $\mathcal{C}$ by 4 and taking the remainder, we obtain the ground truth pixel correspondences $\{\hat{\mathcal{C}}\}$ in each $\mathbf{p}$ matrix, which leads to the second part of our loss function, $\mathcal{L}_{\text{pixel}}$.
@@
where $\mathbf{p}^{i}$ denotes the assignment matrix computed for the $i$-th pair of patches in the matching module. $\hat{\mathcal{C}_{i}}$ represents the $i$-th ground truth pixel correspondence, and $\mathbf{p}^{i}_{\hat{\mathcal{C}_{i}}}$ refers to the value at the corresponding position in $\mathbf{p}^{i}$.
During training, we supervise both parts jointly, resulting in a total loss function that is the sum of the two components, expressed as:
@@
4 Experimental Results and Analysis
Table 1: Registration accuracy on the KITTI and nuScenes datasets. Lower is better for RTE and RRE, higher is better for Acc.
@@ 4 Experimental Results and Analysis
Ours 0.21 ± 0.25 0.67 ± 0.80 99.03 0.82 ± 0.74 0.87 ± 2.12 91.50
4.1 Dataset
As in [39], we conducted experiments on the KITTI Odometry and nuScenes datasets. KITTI Odometry[11]. We generated image-point cloud pairs from the same data frame of 2D/3D sensors. Following previous works, we used sequences 00-08 for training and sequences 09-10 for testing. We downsample the camera image resolution to 512×160 and employ reflectance maps with a resolution of 1024×64 for both training and testing. Artificial errors applied to the point clouds during training and testing include ±10m two-dimensional translations along the x and y axes and arbitrary rotations around the z-axis. It is worth noting that, since the center position in the spherical coordinates of point clouds is typically fixed, translations along the x and y axes applied in other works do not accurately reflect real-world scenarios. To ensure fairness and realism, we initially applied arbitrary rotations around the z-axis to the point cloud before generating reflectance maps. then we applied translations along the x and y axes to the point cloud.
@@ 4.1 Dataset
4.2 Compared Methods
@@ 4.3 Registration Accuracy
where $\gamma(i),i=1,2,3$ are the Euler angles of the matrix $\mathbf{R}_{gt}^{-1}\mathbf{R}_{E}$. Here, the rotation matrix $\mathbf{R}_{gt}$ and the translation vector $\mathbf{t}_{gt}$ denote the transformation of the ground truth, while matrix $\mathbf{R}_{E}$ and vector $\mathbf{t}_{E}$ define the estimated transformation. We calculated the mean and standard deviation of all the data without removing the image-point cloud samples with large errors, as done in CorrI2P. Furthermore, for the registration success rate (Acc.), registrations with $\text{RTE}<2m$ and $\text{RRE}<5$° are considered successful. The final results are presented in Table 1.
Figure 6: Histograms of distributions of RTE and RRE on the KITTI dataset. x-axis is RTE(m) and RRE(°), and y-axis is the percentage. (a) Histogram for proposed method (b) Histogram for the VP2P-Match method (c) Histogram for the CorrI2P method
At the same time, we compared the distribution of RTE and RRE in different intervals with other methods using histograms in Figure 6. We conducted the comparison on the KITTI dataset. A higher proportion in the leftward intervals indicates smaller errors. As can be seen, our method achieves around 90% of matching cases where $\text{RTE}<0.5m$ and $\text{RRE}<1$°, significantly surpassing the other methods, which demonstrates the accuracy of the method we proposed.
4.4 Efficiency Comparison
Table 2: The efficiency comparisons with other methods.
@@ 4.4 Efficiency Comparison
Ours 36.09 0.12
We compared the model size and inference time of our method with other methods on a platform equipped with an NVIDIA GeForce RTX 3090 GPU and an AMD Ryzen Threadripper 2990WX. The results are shown in Table 2. We evaluated our conducted efficiency test on the KITTI dataset, calculating the inference time per frame by dividing the total inference time with the total number of input frames. From the results of the efficiency comparison, we can see that our model has a relatively low number of parameters and short runtime, allowing the registration process to be real-time.
4.5 Visualization Result
We visualized the matching results in Figure 7. Specifically, we present the matching results between the camera image and the reflectance map. The estimated matching pixel pairs are connected by lines in this graph, which will be used for subsequent pose estimation. In the pose estimation process, outlier pairs discarded by RANSAC are marked with red lines, indicating inaccurate estimates, and the accurate matching pairs are connected by green lines. These images show that the estimated matching pairs are not only sufficiently numerous but also highly accurate. Notably, the second figure shows that even when the overlapping region of the point cloud and the image are projected onto the boundary area, we can still estimate the correspondences well, which prove the robustness of our method.
@@ 4.6 Ablation Study
First, we primarily conduct ablation experiments on the range map and reflectance map branches in the feature extraction network to prove the effectiveness of the additional information. The results are shown in Table 3. This table shows that neither branch achieves the best accuracy without the additional information. In some scenarios where point cloud degradation occurs within the camera’s field of view, relying solely on the range map’s point cloud geometric information fails to provide good registration. The additional reflectance information helps to address the issues in such scenes. Similarly, if only the reflectance map is used without the geometric information for assistance, it cannot handle some special cases where the reflectance information is not enough.
Second, we tested the impact of different selections of top-k value during the patch matching stage on registration accuracy and inference time, As shown in Table 4. When the top-k value is small, the accuracy slightly decreases, but the computation speed significantly improves. However, increasing the top-k value does not always lead to better performance. In fact, when the top-k value is too large, there is no significant improvement in accuracy, it results in a substantial decrease in efficiency. Therefore, in practical use, it is important to choose an appropriate value that balances both accuracy and speed.
Table 3: Experimental results on ablation of feature extraction network branches
@@
Table 4: Results of ablation experiments on top-k selection. ‘Time’ for the inference time.
@@
5 Conclusions
We propose a framework for cross-modal registration tasks by projecting 3D LiDAR point clouds into 2D representations. The projection method better leverages the geometric properties of the point clouds while reducing the impact of cross-modal differences. Furthermore, to address potential issues in the image matching process of the registration task, we introduce a multi-scale feature extraction network and a patch-to-pixel matching strategy to get better adaption. Experiments on the KITTI and nuScenes datasets demonstrate that our network achieves exceptional accuracy, precision, and robustness.

### v2

@@ III-E Loss Function
During training, we do not use top-k in the matching network, instead, we use the ground truth patch correspondences for subsequent pixel-level matching directly. For each patch correspondence, the network calculates $M$ assignment matrices $\mathbf{p}$. By dividing the actual correspondences $\mathcal{C}$ by 4 and taking the remainder, we obtain the ground truth pixel correspondences $\{\hat{\mathcal{C}}\}$ in each $\mathbf{p}$ matrix, which leads to the second part of our loss function, $\mathcal{L}_{\text{pixel}}$:
@@
where $\mathbf{p}^{i}$ denotes the assignment matrix computed for the $i$-th pair of patches in the matching module, $\hat{\mathcal{C}_{i}}$ represents the $i$-th ground truth pixel correspondence, and $\mathbf{p}^{i}_{\hat{\mathcal{C}_{i}}}$ refers to the value at the corresponding position in $\mathbf{p}^{i}$.
During training, we supervise both parts jointly, resulting in a total loss function that is the sum of the two components, expressed as
@@
IV Experimental Results and Analysis
TABLE I: Registration accuracy on the KITTI and nuScenes datasets. Lower is better for RTE and RRE, higher is better for Acc.
@@ IV Experimental Results and Analysis
EEPNet[17] 0.54 ± 0.85 2.97 ± 3.05 85.74 - - -
EEPNet-V2 0.21 ± 0.25 0.67 ± 0.80 99.03 0.82 ± 0.74 0.87 ± 2.12 91.50
    * The vacant data in the table indicates that the method could not be effectively trained on this dataset.
IV-A Dataset
We conducted experiments on the KITTI Odometry and nuScenes datasets.
KITTI Odometry[46]. We generated image-point cloud pairs from the same data frame of 2D/3D sensors. Following previous works, we used sequences 00-08 for training and sequences 09-10 for testing. We downsample the camera image resolution to 512×160 and employ reflectance maps with a resolution of 1024×64 for both training and testing. Artificial errors applied to the point clouds during training and testing include ±10m two-dimensional translations along the x and y axes and arbitrary rotations around the z-axis. It is worth noting that, since the center position in the spherical coordinates of point clouds is typically fixed, translations along the x and y axes applied in other works do not accurately reflect real-world scenarios. To ensure fairness and realism, we initially applied arbitrary rotations around the z-axis to the point cloud before generating reflectance maps. then we applied translations along the x and y axes to the point cloud.
@@ IV-A Dataset
Figure 7: Histograms of distributions of RTE and RRE on the KITTI dataset. x-axis is RTE(m) and RRE(°), and y-axis is the percentage. (a) Histogram for RTE (b) Histogram for RRE
IV-B Compared Methods
@@ IV-C Registration Accuracy
where $\gamma(i),i=1,2,3$ are the Euler angles of the matrix $\mathbf{R}_{gt}^{-1}\mathbf{R}_{E}$. Here, the rotation matrix $\mathbf{R}_{gt}$ and the translation vector $\mathbf{t}_{gt}$ denote the transformation of the ground truth, while matrix $\mathbf{R}_{E}$ and vector $\mathbf{t}_{E}$ define the estimated transformation. We calculated the mean and standard deviation of all the data without removing the image-point cloud samples with large errors, as done in CorrI2P. Furthermore, for the registration success rate (Acc.), registrations with $\text{RTE}<2m$ and $\text{RRE}<5$° are considered successful. The final results are presented in Table I.
At the same time, we compared the distribution of RTE and RRE in different intervals with other methods using histograms in Figure 7. We conducted the comparison on the KITTI dataset. A higher proportion in the leftward intervals indicates smaller errors. As can be seen, our method achieves around 90% of matching cases where $\text{RTE}<0.5m$ and $\text{RRE}<1$°, significantly surpassing the other methods, which demonstrates the accuracy of the proposed method.
IV-D Efficiency Comparison
TABLE II: The efficiency comparisons with other methods.
@@ IV-D Efficiency Comparison
EEPNet 58.84 0.015
EEPNet-V2 36.09 0.12
We compared the model size and inference time of our method with other methods on a platform equipped with an NVIDIA GeForce RTX 3090 GPU and an AMD Ryzen Threadripper 2990WX. The results are shown in Table II. We evaluated our conducted efficiency test on the KITTI dataset, calculating the inference time per frame by dividing the total inference time with the total number of input frames. From the results of the efficiency comparison, we can see that our model has a relatively low number of parameters and short runtime, allowing the registration process to be real-time.
IV-E Visualization Result
Fig. 8 presents a comparison of the registration results between different methods on the KITTI dataset, displaying the projections of the point clouds in the image plane. The color of the points in the images represents the distance from the camera. We show the initial unaligned projection results (Input), the aligned projection results obtained using the ground truth (GT), and the projection results after registration using different registration methods. It shows that our method achieves the the most accurate registration result in different environments.
Besides, we visualized the matching results in Fig. 9. Specifically, we present the matching results between the camera image and the reflectance map. The estimated matching pixel pairs are connected by lines in this graph, which will be used for subsequent pose estimation. In the pose estimation process, outlier pairs discarded by RANSAC are marked with red lines, indicating inaccurate estimates, and the accurate matching pairs are connected by green lines. These images show that the estimated matching pairs are not only sufficiently numerous but also highly accurate. Notably, the second figure shows that even when the overlapping region of the point cloud and the image are projected onto the boundary area, the proposed method can still estimate the correspondences well, which prove the robustness of our method.
Figure 8: Visualization of registration results, demonstrating the projection of LiDAR point clouds onto the camera plane after transformation using different extrinsic calibration methods. The color of projected points indicates their distance from the sensor (range), with warmer hues representing closer points and cooler hues representing farther points.
@@ IV-F Ablation Study
First, we primarily conduct ablation experiments on the range map and reflectance map branches in the feature extraction network to prove the effectiveness of the additional information. The results are shown in Table III. This table shows that neither branch achieves the best accuracy without the additional information. In some scenarios where point cloud degradation occurs within the camera’s field of view, relying solely on the range map’s point cloud geometric information fails to provide good registration. The additional reflectance information helps to address the issues in such scenes. Similarly, if only the reflectance map is used without the geometric information for assistance, it cannot handle some special cases where the reflectance information is not enough.
Second, we tested the impact of different selections of top-k value during the patch matching stage on registration accuracy and inference time, As shown in Table IV. When the top-k value is small, the accuracy slightly decreases, but the computation speed significantly improves. However, increasing the top-k value does not always lead to better performance. In fact, when the top-k value is too large, there is no significant improvement in accuracy, it results in a substantial decrease in efficiency. Therefore, in practical use, it is important to choose an appropriate value that balances both accuracy and speed.
TABLE III: Experimental results on ablation of feature extraction network branches
@@
TABLE IV: Results of ablation experiments on top-k selection. ‘Time’ for the inference time.
@@
V Conclusion
We propose a framework for cross-modal registration task by projecting 3D LiDAR point clouds into 2D representations. The projection method better leverages the geometric properties of the point clouds while reducing the impact of cross-modal differences. Furthermore, to address the limitation in the previous method EEPNet in the image matching process of the registration task, we introduce a multi-scale feature extraction network and a patch-to-pixel matching strategy to get better adaption. Experiments on the KITTI and nuScenes datasets demonstrate that our network achieves exceptional accuracy, precision, and robustness.